OpenStudy (anonymous):

MEDAL What are the first 4 steps for graphing this quadratic function? y = -x^2 - 4x - 3

OpenStudy (anonymous):

not sure what specific answers are you looking but here's what i'm gonna do. 1. find the x-intercepts/y-int by factoring the equation 2. Plot them to the xy plane 3. find the point of the vertex (max/min point) and plot it 4. Connect the dots in a smooth curve

OpenStudy (jdoe0001):

hmmm I'd think pick a few points and graph OR find the x-intercepts by getting its roots keep in mind that \(\bf y = -x^2 - 4x - 3\implies y=-({\color{brown}{ x^2+4x+3}}) \) so set y = 0 and factor \(\bf 0=-({\color{brown}{ x^2+4x+3}})\) from there you can find its vertex